OpenStudy (anonymous):

Solve the polynomial inequality. Express the solution set in interval notation. (x/x-3) > 4

OpenStudy (anonymous):

x<4

OpenStudy (anonymous):

interval notation?

OpenStudy (anonymous):

is it asking for a solution set?

OpenStudy (anonymous):

you have to write \[\frac{x}{x-3}>4\] then \[\frac{x}{x-3}-4>0\] then \[\frac{x-4(x-3)}{x-3}>0\] and \[\frac{-3x+12}{x-3}>0\] and solve that one

OpenStudy (anonymous):

thank you againnn!